OpenStudy (anonymous):

A rectangular playground is enclosed on three sides by a fence and one side of a house. The length of the fence is 20 yards. Given that the width of the playground is x yards, find: an expression for the lenth og the playground in temrs of x the area of the playground in terms of x the actual length and width of the playground that will give the maximum area this is a caculus question for optimization so...yeah

OpenStudy (anonymous):

\[l=20-2x\]

OpenStudy (anonymous):

how do you know that?

OpenStudy (anonymous):

\[A=x(20-2x)=20x-2x^2\]\]

OpenStudy (anonymous):

how do you know that...?

OpenStudy (anonymous):

i know it because the total is 20.

OpenStudy (anonymous):

yeah...

OpenStudy (anonymous):

you have 20 yards and one side is against the house. if we call this L then \[l+2x=20\] and so \[l=20-2x\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you only have to use up 3 sides for your 20 because one is free (against the house)

OpenStudy (anonymous):

and width is 2x and area is l*w so A=2(20-x)...?

OpenStudy (anonymous):

normally perimeter of a rectangle would be \[2l+2w\] but in this case one side is free so it is \[l+2x=20\]

OpenStudy (anonymous):

area is width times length so next answer is \[A=x(20-2x)\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

or \[A(x)=20x-2x^2\]

OpenStudy (anonymous):

now you sure as heck don't need calculus to find the maximum area because this is a parabola that faces down and its maximum is at the vertex.

OpenStudy (anonymous):

if you recall that the vertex is given by \[-\frac{b}{2a}\] you are done.

OpenStudy (anonymous):

dont we find the derivative and then solve for x?

OpenStudy (anonymous):

but since you are supposed to use calculus take the derivative, set = 0 and solve. you will still get \[-\frac{b}{2a}\]

OpenStudy (anonymous):

you mean "take the derivative, set = 0 and solve for x" yes

OpenStudy (anonymous):

yes...

OpenStudy (anonymous):

\[A'(x)=20-4x\]

OpenStudy (anonymous):

set = 0 get \[20-4x=0\] \[x=5\]

OpenStudy (anonymous):

but i repeat: you do not need calculus for this. the vertex is \[-\frac{b}{2a}\] whether you take the derivative, set = 0 and solve, or just remember that the vertex of a parabola is given by that.

OpenStudy (anonymous):

in any case x = 5 and the other side is 5 and the side against the house it 10

OpenStudy (anonymous):

okay lol thanks for the tip...-b/2a

OpenStudy (anonymous):

now i you want to impress your math teacher... use P instead of 20

OpenStudy (anonymous):

you will see that in this case the side against the house should be \[\frac{p}{2}\] and the other two sides should be \[\frac{P}{4}\]

OpenStudy (anonymous):

its a practice question for finals...not for a grade but okay let me hear it...lol

OpenStudy (anonymous):

and then do it without the side against the house being free and you will see that each side should be \[\frac{P}{4}\] i.e. a square.

OpenStudy (anonymous):

btw \[-\frac{b}{2a}\] is for \[f(x)=ax^2+bx+c\] for higher degrees you will need calculus

OpenStudy (anonymous):

okay..

OpenStudy (anonymous):

im trying to understand what you said about the P/4 and P/2 i'll just work it out

OpenStudy (anonymous):

thanks so much satellite! :)